I have the graph for the problem, i just need help on figuring out how they got the answer. The equation is (x-2)/3 - 5/(2y-4) =0.
Answer by ewatrrr(24785)

(x-2)/3 - 5/(2y-4) =0
(2y-4)(x-2) - 15 = 0 |Multiplying thru by 3(2y-4) so as all denominators = 1
2yx - 4x - 4y + 8 - 15 = 0
2yx - 4x - 4y = 7
2yx - 4y - 4x = 7
2y(x-2) - 4x = 7
y = (4x+7)/2(x-2)